# Optimal In-Network Distribution of Learning Functions for a Secure-by-Design Programmable Data Plane of Next-Generation Networks

### v1

@@ IV-B Meta-heuristic
IV-C Strong Learner Splitting and #colors Selection
@@ V-B Network Evaluation Campaigns
We evaluated the proposal scalability under three topologies of increasing dimensions: the first one with 10 nodes and 25 edges, for which the value of #$\#colors$ computed by Eq.12 is $3$ (i.e., SL is splitted into three WLs); the second one with 25 nodes and 48 edges and a computed #$\#colors$ equal to $5$ ; and the third, bigger, topology with 30 nodes and 51 edges, for which #$\#colors=7$. The 15-node topology previously considered is not used for these experiments as the calculated value of #$\#colors$ was found to be identical to that of the 10-node topology, and thus it adds little to the experiment.
@@
The proposal has been implemented using P4-enabled virtual PDP, namely BMv2 [33] that are based on the v1Model architecture. Due to the limited instruction set of the P4 language (it does not support basic operations such as division, exponentiation or logarithm), we extracted 43 features of the CIC-IDS 2018. The P4 code that implements the models and the associated feature extractor will be publicly made available. [^2] GitHub repository at [32]
@@ V-B4 Shortest Path Detouring Analysis
Table III presents the average AWDelay values grouped by the number of nodes $N$ and the density class $ED$. The AWDelay values shown for each combination of $N$ and $ED$ represent the average computed across all instances discussed in Section V-A2. The AVG row reports the average calculated based on the nodes, while the column AVG shows the average relative to the density. Additionally, the row labeled VAR indicates the variance of all AWDelay values for each number of nodes $N$, providing a measure of data dispersion and allowing us to assess the variability with the number of nodes.
@@
The results of the experiments, as shown in the Table, indicate that the average weighted delay behaves consistently as the network grows in size. Specifically, AWDelay significantly decreases with an increasing number of nodes. For example, in networks with 10 nodes in the density class $ED1$, the average delay reaches around 5%, while for networks with 30 nodes, the delay drops to approximately 0.5%. This trend can also be observed in the average delay, which decreases from 6.58% with 10 nodes to 0.57% with 30 nodes. This indicates that the overhead introduced by the coloring constraints becomes less significant in larger networks, making the approach more scalable and efficient as the network grows.
@@
TABLE III: AWDelay results for density and number of nodes

### v2

@@ IV-D Meta-heuristic
IV-E Strong Learner Splitting and #colors Selection
@@ V-B Network Evaluation Campaigns
We evaluated the proposal scalability under three topologies of increasing dimensions: the first one with 10 nodes and 25 edges, for which the value of #$\#colors$ computed by Eq.16 is $3$ (i.e., SL is splitted into three WLs); the second one with 25 nodes and 48 edges and a computed #$\#colors$ equal to $5$ ; and the third, bigger, topology with 30 nodes and 51 edges, for which #$\#colors=7$. For the sake of the experimental campaign, we chose as an SL a Random Forest (RF) and, therefore, the WLs are represented by the Decision Trees (DTs) that compose the RF. This choice does not affect the general applicability of the proposed Split-AI technique. Any type of ensemble learning model can be employed, with the only constraint being the instruction set supported by the target programmable network element to encode the model within it. The performance of each considered SL are summarized in Tab.III. As previously mentioned, #$\#colors$ correspond to the number of WLs in the SL.
TABLE III: SL Performance Analysis
Accuracy FPR FNR F1-Score
SL3 $\sim$981.4% 0.15% $\sim$97%
SL5 $\sim$991.3% 0.5% $\sim$98%
SL7 $\sim$990.2% 0.3% $\sim$99%
The 15-node topology previously considered is not used for these experiments as the calculated value of #$\#colors$ was found to be identical to that of the 10-node topology, and thus it adds little to the experiment.
@@
The proposal has been implemented using P4-enabled virtual PDP, namely BMv2 [32] that are based on the v1Model architecture. Due to the limited instruction set of the P4 language (it does not support basic operations such as division, exponentiation or logarithm), we extracted 43 features of the CIC-IDS 2018. The P4 code that implements the models and the associated feature extractor will be publicly made available. [^2] The code will be made available on GitHub repository at [31].
@@ V-B4 Shortest Path Detouring Analysis
Table IV presents the average AWDelay values grouped by the number of nodes $N$ and the density class $ED$. The AWDelay values shown for each combination of $N$ and $ED$ represent the average computed across all instances discussed in Section V-A2. The AVG row reports the average calculated based on the nodes, while the column AVG shows the average relative to the density. Additionally, the row labeled VAR indicates the variance of all AWDelay values for each number of nodes $N$, providing a measure of data dispersion and allowing us to assess the variability with the number of nodes.
@@
The results of the experiments, as shown in Table IV, indicate that the average weighted delay behaves consistently as the network grows in size. Specifically, AWDelay significantly decreases with an increasing number of nodes. For example, in networks with 10 nodes in the density class $ED1$, the average delay reaches around 5%, while for networks with 30 nodes, the delay drops to approximately 0.5%. This trend can also be observed in the average delay, which decreases from 6.58% with 10 nodes to 0.57% with 30 nodes. This indicates that the overhead introduced by the coloring constraints becomes less significant in larger networks, making the approach more scalable and efficient as the network grows.
@@
TABLE IV: AWDelay results for density and number of nodes